# Towards Open-World Retrieval-Augmented Generation on Knowledge Graph: A Multi-Agent Collaboration Framework

### v1

@@ 4.7 Effect of the Retrieval
To further validate the effectiveness of our method in anchor identifying and knowledge retrieval, we compare the effect of retrieval of our method against two RAG baselines (PoG and ToG) and a variant (AnchorRAG w/o Parallel Retrieval). The results are shown in Fig. 3. The figures in the top row illustrate the accuracy of anchor identifying (marked in green), where our method significantly outperforms the baselines by grounding the entities with their textual and neighborhood information. The bottom row presents the proportion of the exact match for the answers, including the answers obtained by retrieval (marked in red) and those inferred by CoT reasoning (marked in orange). Notably, AnchorRAG achieves the highest retrieval rate at 49.6%, indicating that the multi-agent collaborative mechanism can effectively obtain the correct answer entities, which supports the following reasoning process to improve the QA performance.
4.8 Hyper-parameter Analyses
@@
Table 4: The results of AnchorRAG with different hyper-parameters using GPT-4o-mini on GrailQA.
@@
We conducted the parameter sensitivity analyses on Grail-QA to investigate the optimal configuration for the number of neighborhood relations $k$ and retrieval agents $m$, where the results are shown in Table 4. We can find that retaining excessive neighboring relations $k$ for each candidate entity will lead to a performance loss, due to the fact that it may introduce noise and neglect the key structural information. Besides, a proper number of retriever agent $m$ can make AnchorRAG achieve the optimal results, which can reflect the effectiveness of the MAC framework. In addition, we can also find that more retrieval agents won’t always obtain the better results. Excessive agents will introduce uncertain retrieval information into the collaborative process, thereby impairing the performance.
4.9 Case Study
To intuitively demonstrate the superiority of our method in knowledge graph retrieval, we conduct a case study that compares our approach with the ToG baseline on a representative complex question from the CWQ dataset: “Who inspired F. Scott Fitzgerald, and who was the architect that designed The Mount?”. As shown in Fig. 4, we can observe that ToG can’t locate both the anchor entities, resulting in meaningless reasoning paths and generating an incorrect answer. In contrast, our multi-agent framework can effectively identify the anchor entities for the given questions by the entity grounding mechanism. With assigning multi retriever agents to independently explore different reasoning paths, AnchorRAG accurately locates the correct answer entity “Edith Wharton”. Additionally, this case highlights the robustness of our method. For example, although it initially considers an incorrect entity (“The Weapon”) as an anchor entity, it promptly detects the irrelevance of the following retrieval path with the rough pruning and fine filtering method, and terminates the corresponding agent, which prevents unnecessary computation overhead.
Figure 4: Comparisons of retrieval process between a baseline reasoning (ToG) and our method (AnchorRAG) on a complex question. The entities marked in yellow are anchor entities, marked in green are correct answer entities. The ones marked in blue and striped are invalid entities.
5 Conclusion
@@
In this paper, we present AnchorRAG, a novel multi-agent collaborative RAG framework for open-world Knowledge Graph Question Answering. Unlike existing methods that rely on predefined anchor entities, AnchorRAG overcomes the bottleneck of unreliable entity linking by identifying and grounding plausible anchors without prior assumptions. Specifically, a predictor agent dynamically identifies candidate anchor entities by aligning the given questions with the candidate entities and initializes independent retriever agents to conduct parallel multi-hop explorations. Then a supervisor agent formulates the iterative retrieval strategy for these retriever agents and synthesizes the reasoning paths to generate the final answer. Extensive experiments on four public KGQA benchmarks demonstrate that AnchorRAG significantly outperforms the state-of-the-art baselines, validating the effectiveness of our MAC framework, especially when dealing with the complex multi-hop reasoning tasks.

### v2

@@ 5.5. Effect of the Retrieval
To further validate the effectiveness of our method in anchor identifying and knowledge retrieval, we compare the effect of retrieval of our method against two RAG baselines (PoG and ToG) and a variant (AnchorRAG w/o Parallel Retrieval). The results are shown in Figure 4. The figures in the top row illustrate the accuracy of anchor identifying (marked in green), where our method significantly outperforms the baselines by disambiguating the entities with their textual and neighborhood information. The bottom row presents the proportion of the exact match for the answers, including the answers obtained by retrieval (marked in red) and those inferred by CoT reasoning (marked in orange). Notably, AnchorRAG achieves the highest retrieval rate at 49.6%, indicating that the multi-agent collaborative mechanism can effectively obtain the correct answer entities, which supports the following reasoning process to improve the QA performance.
Figure 4. The effect of the anchor identifying and knowledge retrieval on WebQSP.
5.6. Hyper-parameter Analyses
We conducted the parameter sensitivity analyses on GrailQA to investigate the optimal configuration for the number of neighborhood relations $k$ and retrieval agents $m$, where the results are shown in Table 4. We can find that retaining excessive neighboring relations $k$ for each candidate entity will lead to a performance loss, due to the fact that it may introduce noise and neglect the key structural information.
@@ 5.6. Hyper-parameter Analyses
Table 4. The results of AnchorRAG with different hyper-parameters using GPT-4o-mini on GrailQA.
@@
Besides, a proper number of retriever agent $m$ can make AnchorRAG achieve the optimal results, which can reflect the effectiveness of the MAC framework. In addition, we can also find that more retrieval agents won’t always obtain the better results. Excessive agents will introduce uncertain retrieval information into the collaborative process, thereby impairing the performance. Furthermore, we also conduct the hyperparameter sensitivity analysis for retrieval depth and width. The detailed experimental results and analyses can be found in the appendix B.
6. Conclusion
@@
In this paper, we present AnchorRAG, a novel multi-agent collaborative RAG framework for open-world Knowledge Graph Question Answering. Unlike existing methods that rely on predefined anchor entities, AnchorRAG overcomes the bottleneck of unreliable entity linking by identifying and grounding plausible anchors without prior assumptions. Specifically, a predictor agent dynamically identifies candidate anchor entities by aligning the given questions with the candidate entities and initializes independent retriever agents to conduct parallel multi-hop explorations. Then a supervisor agent formulates iterative retrieval strategy for these retriever agents and synthesizes the reasoning paths to generate the final answer. Extensive experiments on four public KGQA benchmarks demonstrate that AnchorRAG significantly outperforms the state-of-the-art baselines, validating the effectiveness of our MAC framework, especially when dealing with the complex multi-hop reasoning tasks.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (62476274, U22B2048, 62394330), in part by the Special Project of Joint Funding for Municipal, University (Institute) and Enterprise under the Guangzhou Basic Research Program (2024A03J0395), in part by the Science and Technology Projects in Guangzhou (2024D03J0010).
Appendix A Details of Experimental Datasets
We use four benchmark datasets: WebQuestionSP (WebQSP) (Yih et al., 2016), GrailQA (Gu et al., 2021) and Complex WebQuestions (CWQ) (Talmor and Berant, 2018), and WebQuestions (WebQ) (Berant et al., 2013) to evaluate the effectiveness of our method. These datasets include a training set, a validation set, and a test set. For fair comparison, we follow the baselines (Sun et al., 2023; Chen et al., 2024) and use only the test sets as evaluation benchmarks.
Datasets #Train #Validation #Test
WebQSP 3098 - 1639
GrailQA 44,337 6763 13231
CWQ 27734 3480 3475
Webquestions 3778 - 2032
WebQSP(Open-World) 3098 - 1803
GrailQA(Open-World) 44,337 6763 1100
Table 5. Statistics of datasets.
Datasets 1 hop 2 hop $\geq$ 3 hop
WebQSP 65.49% 34.51% 0.00%
GrailQA 68.92% 25.82% 5.26%
CWQ 40.91% 38.34% 20.75%
Table 6. statistics of the reasoning hops in WebQSP, CWQ and GrailQA.
In order to save computational resources, following prior research (Sun et al., 2023; Chen et al., 2024), we only used a sample of 1000 data points from the GrailQA dataset for evaluation. Noticing that, in order to verify the performance of each method in open world settings, we introduce noise into the WebQSP and GrailQA datasets, rewriting keywords in the questions to make it impossible to directly match entities in the knowledge graph through string matching. We also sample a small portion of questions from two other datasets, HotpotQA (Yang et al., 2018) and TriviaQA (Joshi et al., 2017), and merge them with the noisy data from WebQSP and GrailQA to create two open world datasets (Open-World WebQSP and Open-World GrailQA) to verify the robustness of our method. The statistics of the datasets are shown in Table 5, and the statistics of the reasoning hops are shown in Table 6.
Appendix B Experiments of Additional Hyper-parameter
Figure 5 shows the hyperparameter sensitivity analysis for retrieval depth and width. The experimental results strongly validate the effectiveness of our default configuration, i.e., width=3 and depth=3. On both datasets, the Hit@1 performance of our method peaks when these settings are used. This non-linear performance trend is attributed to two reasons. First, the performance improvement is limited by the inherent complexity of the datasets, as most questions do not require more than 3 hops of reasoning, as shown in table 6. Therefore, a deeper search does not yield significant benefits. Second, excessively expanding the search space introduces a significant amount of noise from irrelevant entities and relations, which have the negative impact on the reasoning and lead to a poorer performance. This is shown by the performance degradation on GrailQA when depth or width is increased to 4. In particular, the WebQSP dataset does not contain questions requiring more than 3 hops. Consequently, increasing the retrieval depth beyond 3 on WebQSP offers no performance benefit, and we therefore did not perform experiments with depth=4 on this dataset. To sum up, our parameter selection presents an optimal trade-off between maintaining adequate search space, mitigating noise, and balancing computational overhead.
Figure 5. Performances of AnchorRAG with different search depths and widths.
Appendix C Experiments on Additional LLM Backbones
To further evaluate the universality and robustness of AnchorRAG with different Large Language Model (LLM) backbones, we extend our experiments using two additional state-of-the-art models: Claude4.5-haiku and Gemini-2.5-flash. We compared the performance of our method against the strongest baseline, i.e., ToG, on the GrailQA dataset. The experimental results are presented in Table 7. Consistent with the results in the main experiments, AnchorRAG demonstrates superior performance over the ToG on both new backbones. For example, on the GrailQA dataset, AnchorRAG achieves substantial improvements in Hit@1, surpassing ToG by 13.7% with Claude4.5-haiku and 16.5% with Gemini-2.5-flash. This significant improvement validates that our multi-agent collaboration framework effectively mitigates the entity linking issues in complex open-world scenarios, regardless of the underlying LLM. These findings suggest that AnchorRAG’s effectiveness is model-agnostic and can cooperate with various LLMs to enhance open-world reasoning capabilities.
Dataset GrailQA
Hit@1 Acc F1
Claude4.5-haiku
ToG 31.7 26.6 27.6
AnchorRAG 45.4 39.0 40.3
Gemini-2.5-flash
ToG 39.6 32.8 34.0
AnchorRAG 56.1 47.9 49.7
Table 7. Performance comparison of ToG and AnchorRAG using additional LLM backbones (Claude4.5-haiku and Gemini-2.5-flash) on GrailQA datasets.
Appendix D Case Study
To intuitively demonstrate the superiority of our method in knowledge graph retrieval, we conduct a case study that compares our approach with the ToG baseline on a representative complex question from the CWQ dataset: “Who inspired F. Scott Fitzgerald, and who was the architect that designed The Mount?”. As shown in Figure 6, we can observe that ToG can’t locate both the anchor entities, resulting in meaningless reasoning paths and generating an incorrect answer. In contrast, our multi-agent framework can effectively identify the anchor entities for the given questions by the entity grounding mechanism. With assigning multi retriever agents to independently explore different reasoning paths, AnchorRAG accurately locates the correct answer entity “Edith Wharton”. Additionally, this case highlights the robustness of our method. For example, although it initially considers an incorrect entity (“The Weapon”) as an anchor entity, it promptly detects the irrelevance of the following retrieval path with the rough pruning and fine filtering method, and terminates the corresponding agent, which prevents unnecessary computation overhead.
Figure 6. Comparisons of retrieval process between a baseline reasoning (ToG) and our method (AnchorRAG) on a complex question. The entities marked in yellow are anchor entities, marked in green are correct answer entities. The ones marked in blue and striped are invalid entities.
Appendix E Prompts
The prompts used in our method are detailed in this section.
Prompt for Candidate Entity Generation Extract all topic entities from the given multi-hop question. Topic entities are proper nouns, named entities, or specific concepts that are crucial for retrieving external knowledge. They may come from different sub-questions that contribute to the final answer. If there are multiple topic entities, separate them with commas. In-Context Few-shot Question: <Question>
Prompt for Relations Pruning Please retrieve %s relations (separated by semicolon) that contribute to the question and rate their contribution on a scale from 0 to 1 (the sum of the scores of %s relations is 1). In-Context Few-shot Question: <Question>
Topic Entity: <Topic Entity>
Relations: <Relations>
Prompt for Entities Pruning Please score the entities’ contribution to the question on a scale from 0 to 1 (the sum of the scores of all entities is 1). In-Context Few-shot Question: <Question>
Relation: <Relation>
Entities: <Entities>
Prompt for Triples Filtering Given the question and a list of knowledge triples, identify and return only the triples that are directly relevant to answering the question. Do not change the format of the triples. Do not generate any explanations or extra text. Return only the filtered triples as-is. In-Context Few-shot Question: <Question>
Triples: <Triples>
Prompt for Triples Evaluating Given a question and the associated retrieved knowledge graph triplets (entity, relation, entity), you are asked to answer whether it’s sufficient for you to answer the question with these triplets and your knowledge (Yes or No). In-Context Few-shot Question: <Question>
Triples: <Triples>